Easy Car Corp. is a grocery store located in the Southwest. It plans to pay an annual dividend of ​$2.00 next year to its shareholders and plans to increase the dividend annually at the rate of 3.0​%. It currently has 1,000,000 common shares outstanding. The shares currently sell for ​$12 each. Easy Car Corp. also has 20,000 semiannual bonds outstanding with a coupon rate of 7​%, a maturity of 21 ​years, and a par value of ​$1,000. The bonds currently have a yield to maturity​ (YTM) of 7%. What is the weighted average cost of capital (WACC) for Easy Car Corp. if the corporate tax rate is 20%?

When answering this problem enter your answer using percentage notation but do not use the % symbol and use two decimals (rounding). For example, if your answer is 0.10469 then enter 10.47; if your answer is 10% then enter 10.00
Answer #1

As per Constant Dividend Growth Model,

Stock Price = D1/(r - g)

So,

r = 2/12 + 0.03

r = 19.67%

Cost of Equity = 19.67%

As YTM = 7%

Cost of Debt = 7%\

WACC = WdCd(1 - t) + WeCe

Tax Rate = 20%

Weight of Debt = 20,000,000/(32,000,000) = 0.625

Weight of Equity = 12,000,000/(32,000,000) = 0.375

So,

WACC = 0.625(0.07)(1 - 0.20) + 0.375(0.1967)

WACC = 10.88%